A ball is thrown vertically upward at 32 m/s a. How long will it take for the ball to reach it’s maximum height? b. What will be the ball’s velocity once it returns to the spot where it was thrown?
1
Expert's answer
2020-12-17T09:01:34-0500

How long it will take for the ball to reach its maximum height:


"t=\frac vg=\frac{32}{9.8}=3.27\text{ s}."

When the ball will return to the spot where it was thrown, its velocity will be the same 32 m/s:


"h=\frac{v^2}{2g},\space u=\sqrt{2gh}=v=32\text{ m/s}."